Can we apply pv=mrt to flow process ? if yes how ? if not why?

Grade:12th pass

1 Answers

Shailendra Kumar Sharma
6 years ago
PV= nRT
flow conditions 
d(PV)/dt =d(nRT)/dt
in constant temp and pressure
PdV/dT =RTdn/dT
and you can adjust as your convinience